OpenStudy (anonymous):

A farmer uses pesticides to keep harmful bugs away from his crops. The amount of pesticide used is an inverse variation of the percentage of bugs remaining on the crops. If he uses 5 gallons of pesticide per acre, 5% of the bugs remain. Which equation describes this relationship? (Let p denote the percentage of bugs and g denote the number of gallons used in the equations below.)

OpenStudy (anonymous):

i think it is p=1/4g

OpenStudy (anonymous):

@Owlcoffee @MrNebula @kdeterline

OpenStudy (anonymous):

or if you cant help with that one can you help with this one? If y varies inversely with x and y = 3 when x = 4, what is the value of y when x = 2?

OpenStudy (anonymous):

and the answer choices are A. y=1/6 B. y = 2 C. y = 6 D. y = 12

OpenStudy (mrnebula):

Im on it!

OpenStudy (mrnebula):

Your think your right on the first one

OpenStudy (anonymous):

ya I'm pretty sure i have the first one right

OpenStudy (mrnebula):

For the second one...

OpenStudy (mrnebula):

I think your right on the first one* sorry I had Your two times there xD

OpenStudy (owlcoffee):

So, let's try the equation you came up with (p=1/4g) Replacing with the initial p=5 and g=5 \[p=\frac{ 1 }{ 4 }g\] \[5=\frac{ 1 }{ 4 }(5) \rightarrow 5\neq \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

haha its okay and thats not one of the answer choices for the first one

OpenStudy (mrnebula):

Owl has it down better

OpenStudy (anonymous):

I just really need help on the second one

OpenStudy (anonymous):

nevermind i figured it out. The second one is 6 and the first one is p=1/4g

OpenStudy (anonymous):

Thanks for your help tho cx

OpenStudy (mrnebula):

Here is a example this can help you to learn how to solve other problems like this. If x varies inversely as y and x=12 when y=3, what is the value of x when y=9? y=k/(x) where k is some constant 3=k/12 36=k When y=9: 9=36/x x=4

OpenStudy (mrnebula):

Hope this will help!